A charge ‘Q’ is placed at the center of a cube of side l , what is the electric flux passing through two opposite faces of the cube is
1
Expert's answer
2021-08-05T14:21:36-0400

We know that

Gauss law of electric flux

"\phi=\frac{Q}{\epsilon_0}\rightarrow(1)"

Where Q=charge

equation (1) represent electric flux due to six face of cube

electric flux due to each face of cube

"\phi'=\frac{\phi}{6}\rightarrow(2)"

equation (1) and eqution (2) we can written as

"\phi'=\frac{Q}{6\epsilon}"

the electric flux passing through two opposite faces of the cube

"\phi'=2\times\frac{Q}{6\epsilon_0}=\frac{Q}{3\epsilon_0}"